OpenStudy (anonymous):

HELP ME If g(x) = 3x - 5, then g -1(x)=

OpenStudy (anonymous):

first find the inverse.

OpenStudy (anonymous):

g-1 (x) = ⅓ x + 1 ⅔

OpenStudy (anonymous):

okay then what

OpenStudy (anonymous):

that is the answer. g-1 (x) = ⅓ x + 1 ⅔

OpenStudy (anonymous):

those arent the possible answers so thats incorrect

OpenStudy (anonymous):

that is correct.

OpenStudy (anonymous):

ill get someone to make sure . @mathmale am i right ?

OpenStudy (anonymous):

is it possible that the mixed # should be a improper fraction ?

OpenStudy (mathmale):

Plohrr: I'll be right with you; I'm going to check what you've done. Glad to hear from you.

OpenStudy (mathmale):

Nice work, Plohrr! Would you please explain the steps of the process of finding the inverse of g(x) to peachesandcream? Thanks.

OpenStudy (anonymous):

she said its incorrect !

OpenStudy (mathmale):

It wouldn't matter if you ended up with an improper fraction. There are various ways in which to write the inverse function. Personally, I'd write it f^(-1) = (x+5)/3. But this is exactly equivalent to your result.

OpenStudy (anonymous):

soo. maybe the choice is different in the book

OpenStudy (mathmale):

I'd suggest you suggest the correct result is -1 x + 5 f (x) = ----------- 3 Do you, Plohrr, know how to check whether a function f(x) and its inverse f^(-1) actually are inverses? Hint: f^(-1) (f(x)) MUST EQUAL x.

OpenStudy (mathmale):

I've checked it myself and find that our result absolutely is correct. Again, Plolrr, expressions such as this can be expressed in various ways. Again I assert that y our initial response was correct, just as mine is correct.

OpenStudy (mathmale):

BRB

OpenStudy (anonymous):

ok. thanks !

OpenStudy (mathmale):

Are you satisfied with this discussion? And is your "client" also satisfied?

OpenStudy (anonymous):

i don't know she's away haha. but i definitely understand now.

OpenStudy (anonymous):

see i knew the other thing wasnt right

OpenStudy (anonymous):

k.